Home‎ > ‎Crew 131 - Embry Riddle‎ > ‎

1213-CommandersReport

Commander Report
12/13/2013
Chelsea Iwig

This morning Crew 131 awoke bright and early to complete the high intensity workout for the
exercise study. The crew did an excellent job completing the workout and already seems to be
getting accustomed to it. Following the workout, the crew ate breakfast and completed their daily
preparations. Today was a walking EVA day for team 1, so they headed out to a dried up creek bed
west of the Hab in the late morning. The crew was definitely feeling a little worn down though and
not too enthusiastic about trekking the nearly 2km to the creek bed. However, they pushed forward
and successfully completed their EVA, and the even enjoyed some light hearted games in the snow on
their way back to the Hab.

During the time that team 1 was out on their EVA, team 2 washed dishes and cooked a late
lunch to be ready and on the table when team 1 returned to the Hab. At 1326, team 1 burst through
the airlock door into the Hab and quickly doffed their suits and rushed up the Hab stairs feeling
famished and ready to eat the delicious pesto pasta the team 2 had prepared for them. The crew all
agreed that it had been the best meal they had eaten since arriving on the Red Planet. After lunch,
the crew was in good spirits and ready to begin working on the glove study and other science
projects. However, shortly after lunch, the crew began to have more trouble with the Hab heater
blowing out cold for the second time that afternoon. Meanwhile, the crew engineer also noticed that
the water pump wasn’t working, so she quickly went downstairs to flip the breaker, which resolved
the problem of the heater and pump not working properly. However, in the process of rushing
downstairs to flip the breaker, she had forgotten that the water pump was still on and the indoor
tank ended up running over and water began spilling onto the loft floor. This resulted in a hit to
the previously high crew morale as they now had to mop up the puddle of water that had leaked onto
the loft floor and one crewmember’s bed below. However, the crew cleaned up the water together and
soon the problem was successfully resolved.

After the brief excitement from the water leak, the crew had a brief break to unwind a
little before completing the rest of their research for the day, including usability surveys, glove
testing, and behavioral surveys. Tonight, the crew journalist will be heading out to the Mars
Observatory to get the telescope set up for star observations later tonight. The crew will end the
evening by playing a team communication game, called Space Team, on the iPads to improve moral and
overall team cohesion.

__________________________________


CommanderCheckIn Report

12/13/2013

Chelsea Iwig

Crew Physical Health: One Crew member had some lower back pain after returning from today’s EVA. He will not be going on any more EVAs until further notice. All other crewmembers are in good physical health.

Eva Depart Time:

1127

Eva Return Time:

1321

Narrative Of Field Mission Results: The EVA crew went to a dried up creek bed West of the Hab. The armband GPS were tested again and they performed very well. There are still some kinks that need to be ironed out, but they seem to be a useful tool to add to our space suit design.

EVA team one

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

 

Eva Depart Time:

 

Eva Return Time:

 

Narrative Of Field Mission Results:

Hab Maint: The pipes were thankfully not frozen this morning likely due to our crew engineer turning on the pump every two hours during the night.

ATVs 3 and 4 are still not working. The crew engineer took the batteries out again to charge them, but this did not help. We believe there is something else wrong with them besides the batteries. We will not be able to use the ATVs for EVAs until at least one of them is repaired because we only have two working ATVs.

At 12:45pm the crew noticed that the heater was blowing cool air. I checked the heat vents throughout the HAB and confirmed that no heat was blowing through the HAB. I switched off the breaker for the central air, then opened up the panel. I noticed that 2 wires came out. I re-wired them back to their original position and turned the breaker back on. The heat began to blow through the vents again. Switching the breaker off/on caused power in the HAB to flicker.

At 2:15pm the vents began to blow cool air again and the water pump was not pumping (no sound of water or any other feedback from the pump). I switched the breaker(s) off and on, the pump and the heater reset.

At 2:50pm the heat turned off again. The vents are still blowing but only cool air. All other electrical units are working (including lab heater)

New EVA Suit Testing:

Pack 10: During a walking EVA we found a few issues with the helmet and the pack. The air tubes for the pack do not have the same joints as the old packs, which is creating a large amount of torque on the female pipe that is attached to the helmet. This puts the helmet at risk for cracking (which some of the other new helmets seemed to show). Because of the increased tension, this caused the tubes to pop out of their position (from the pack).

The pack positioning on the frame was higher than the previous packs, which caused lots of discomfort as the weight is not distributed evenly. Despite working on the fans, the air was not blowing nearly as strong as the previous packs. Even if I opened up the bottom, I don't believe this would have increased the fan output.

The pack was incredibly sturdy and a lot more durable than the other packs

The helmet circumference became invasive in regular procedures and was very uncomfortable for the user. There was also a lack of support from the back of the helmet to keep it in place, thus causing the helmet to force the user's chin to their chest. We tried to tighten the tethers at the pack harness of the helmet to the pack frame but it the angle prevented the ability to pull the helmet up very much.

Overall, the user claimed that the pack caused soreness in his back, neck and shoulders due to the positioning and increased weight of the pack.

Also, the material used for the helmet is extremely fragile, and could break very easily on EVAs. It also fogged up even worse than the original helmets. This may also be due to weak airflow from the pack.

Rep Sched:

Commander Check-In, Commander Report, Engineering Report, HSO Report, EVA Report, EVA Plan, Journalist Report, Science Report

Tomorrow Plan:

Tomorrow, the crew will have the opportunity to sleep in a little as the crew will not be exercising in the morning.  Once the crew is up and has finished their daily preparations and breakfast, two crewmembers will conduct the space suit glove study. This will be the first day of the second round of testing for the gloves. All crewmembers have now completed the first round of testing and now the second round will involve different procedures that will test the ability of the crewmembers to complete fine motor tasks while wearing the gloves. Following the glove study, EVA team 2 will go out on their third EVA to a location 2km northeast of the Hab. Following this EVA the crew will continue with the normal afternoon and evening activities, including checking the plants, working on reports, normal Hab maintenance, COMMS, and surveys. In the evening the crew may do some star obs in the observatory or they may stay in the Hab and play team games to boost morale and crew cohesion.

Inventory:

Support Req: New filters for the EVA packs and rat traps are needed. Also, fluorescent light bulbs for the downstairs lab are needed as only one is working currently and it has been flickering on and off.

Miscelleneous: We have not been able to catch any more of the mice or rats as they seem to have become smarter and are now steeling the food off of the traps before the trap springs. We think it may actually just be the rats doing this and we would like to try setting larger rat traps, but we don’t currently have any. Therefore, we are requesting a couple of rat traps to see if they will work better.